OpenStudy (anonymous):

(5x^5y)^3(x/5y)^2 I got 125x^6y^4 but thats wrong and I am really confused

OpenStudy (anonymous):

\[\left( 5x ^{5}y \right)^{3}\left( \frac{ x }{ 5y} \right)^{2}=\frac{ 5^{3}x ^{15}y ^{3}x ^{2 } }{ 5^{2}y ^{2} }\] \[=5^{3}5^{-2}x ^{15+2}y ^{3}y ^{-2}\] \[=5^{3-2}x ^{17}y ^{3-2}=5x ^{17}y\]

OpenStudy (anonymous):

thanks you